### v2

# Can GenAI Move from Individual Use to Collaborative Work? Experiences, Challenges, and Opportunities of Integrating GenAI into Collaborative Newsroom Routines

Abstract.
Generative AI (GenAI) is reshaping work, but adoption remains largely individual and experimental rather than integrated into collaborative routines. Whether GenAI can move from individual use to collaborative work is a critical question for future organizations. Journalism offers a compelling site to examine this shift: individual journalists have already been disrupted by GenAI tools; yet newswork is inherently collaborative relying on shared routines and coordinated workflows. We conducted 27 interviews with newsrooms managers, editors, and front-line journalists in China. We found that journalists frequently used GenAI to support daily tasks, but value alignment was safeguarded mainly through individual discretion. At the organizational level, GenAI use remained disconnected from team workflows, hindered by structural barriers and cultural reluctance to share practices. These findings underscore the gap between individual and collective adoption, pointing to the need for accounting for organizational structures, cultural norms, and workflow integration when designing GenAI for collaborative work.
GenAI, journalism, organizations, teamwork, collaborative work, human-AI collaboration
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Human-centered computing Computer supported cooperative workccs: Human-centered computing Empirical studies in collaborative and social computingccs: Human-centered computing Empirical studies in HCI
1. Introduction
Organizations today face a profound question: How should today’s organizations adopt generative artificial intelligence (GenAI), and how could GenAI become a routine part of collaborative work in future organizations? In 2025, Stanford HAI and Google DeepMind (Stanford Institute for Human-Centered Artificial Intelligence and Google DeepMind, 2025) invited researchers worldwide to explore future collaborative work within organizations, emphasizing that “AI is transforming how we work, make decisions, and collaborate. But while our tools are evolving fast, our organizational models are overdue for reinvention.” Despite its transformative potential, GenAI has entered most workplaces slowly and cautiously. Early adoption often takes the form of individual experiments and tentative trials, rather than sweeping integration led by organizations (Xiao et al., 2025b; Wu, 2024). Across industries, organizations are grappling with a familiar tension in technology adoption: the promise of innovation collides with the uncertainty of organizational change, as leaders and practitioners question how GenAI will reshape collaborative routines and align, or misalign, with teamwork norms (Gartner, 2024). In this study, we investigate how GenAI might move beyond fragmented individual use to become an integral part of organizational collaboration, and what challenges and opportunities this transition creates for reconfiguring collaborative work around GenAI in organizations.
In this paper, we use the newsroom as our primary empirical site to examine what collaborative work with GenAI entails in practice. News production is rarely the effort of a single individual: front-line journalists gather information, draft stories, and exchange leads with peers; their work is then reviewed, edited, and sometimes rewritten by editors; and senior managers provide strategic oversight, shape policies, and liaise with external stakeholders (Duffy, 2021; Xiao et al., 2025b; Tuchman, 1973). Collaboration in newsrooms therefore spans both inter-role and intra-role interactions, anchored in shared routines and collective values such as editorial responsibility and mutual trust (Tuchman, 1973; Carlson, 2009; Robinson, 2011). We distinguish between three key roles around collaborative newswork in our study: front-line journalists, who are primarily responsible for story development and on-the-ground reporting; editors, who oversee content refinement, ensure editorial consistency, and coordinate team workflows; and senior managers, the newsroom leaders responsible for organizational decision-making and long-term strategy. While journalism scholarship often uses the umbrella term journalists to refer broadly to various news practitioners (Duffy, 2021), we adopt this convention as well. In our analysis of how GenAI reshapes collaborative practices and team dynamics, we therefore attend to collaboration within roles as well as across roles.
While these roles structure collaboration in conventional newsroom routines, the rise of GenAI introduces new uncertainties about how such collaboration should unfold. Recently, on the one hand, many individual journalists are experimenting with GenAI in tentative, ad hoc ways (Liang et al., 2025), often relying on personal judgment to ensure that its use remains consistent with professional norms (Wu, 2024). On the other hand, major news organizations have started to articulate institutional responses to the rise of GenAI, for example, China Media Group (CMG) is developing proprietary GenAI systems for its news production (China Press, Publication, Radio, Film and Television Journal, 2024; Lang, 2024), while The British Broadcasting Corporation (BBC) has issued formal usage guidelines for individual journalists (BBC News, 2024). However, these responses have not yet resolved how GenAI should be integrated into collaborative newsroom practices: how teams of journalists and editors might coordinate GenAI use and maintain shared editorial standards in future collaborations.
To understand why this missing layer of collaboration matters, we turn to the normative features of journalism that make it a particularly high-stakes and revealing site for studying GenAI’s integration into collaborative work. First, as a content-intensive profession, journalism is especially exposed to GenAI’s disruptions: core newsroom tasks such as brainstorming, reporting, and editing directly overlap with GenAI’s generative capabilities in text production, compelling journalists to negotiate how these systems intervene in their everyday collaborations (Cools and Diakopoulos, 2024). Second, newswork is inherently collaborative, structured through collective editorial routines and shared news values (Schudson, 2001; Deuze, 2005). GenAI therefore does not simply alter individual tasks but also potentially reshapes the collaborations and mutual trust on which newsroom collaboration depends. Third, the public visibility of journalism subjects GenAI adoption to heightened scrutiny, both inside and outside the newsroom. This visibility amplifies the stakes of collaborative decisions about when and how to use GenAI, since any misstep is not only organizationally consequential but also publicly accountable (Petre, 2021; Chadha and Koliska, 2016; Nishal and Diakopoulos, 2025; Møller, 2024; Komatsu et al., 2020).
Yet despite journalism’s collaborative nature, much of the existing research on GenAI use has remained focused at the individual level. Early research on GenAI in newsrooms has largely focused on task-level affordances and risks for individuals. Scholars have highlighted how AI tools can accelerate content creation and support editorial work, while raising concerns about factual errors, bias amplification, content homogenization, and the erosion of core journalistic skills (Noh et al., 2025; Wang et al., 2025; Woodruff et al., 2024; Siitonen et al., 2024). More recent work has turned attention to value alignment and professional norms, showing that individual journalists adopt GenAI selectively and cautiously, favoring uses that preserve editorial credibility (Nishal and Diakopoulos, 2025; Tseng et al., 2025; Xiao et al., 2025b; Wu, 2024). However, a key gap remains: we know little about how GenAI becomes, or fails to become, a part of collaborative newsroom routines. By this we mean not sporadic or experimental uses of GenAI, but its sustained integration into everyday collaborative newswork, where its presence is taken for granted, embedded in daily teamwork, and aligned with collaborative norms.
Therefore, our study asks:
  • RQ1: What are journalists’ current experiences with GenAI that intersect with collaborative newsroom routines?
  • RQ2: What challenges do journalists face in integrating GenAI into their collaborative newsroom routines?
  • RQ3: How can newsrooms better integrate GenAI into collaborative newsroom routines in the future?
To answer these questions, we conducted an in-depth qualitative study of Chinese newsrooms. China is a fertile site for studying GenAI adoption in newsroom collaborations because Chinese newsrooms operate under both rapid technological transformation and strict professional expectations. Journalists navigate a dual pressure: leveraging cutting-edge GenAI tools to enhance innovation under governmental expectations, while upholding the editorial authority that underpins their professional norms (Xiao et al., 2025b). This dual pressure magnifies the challenges of integrating GenAI into collaborative newsroom routines and makes teamwork negotiation around GenAI especially visible.
Our study draws on 27 semi-structured interviews with senior newsroom managers, editors, and front-line journalists in China who had participated in early GenAI adoption in their newsrooms. Our findings showed that GenAI is already widely used at the individual level in Chinese newsrooms, but such use remains highly personalized, informal, and largely disconnected from collaborative routines (RQ1). Then, we found that structural fragmentation at the organizational level and cultural reluctance to share practices have prevented GenAI from being integrated into collaborative newsroom routines (RQ2). Journalists also envision GenAI-augmented collaboration as requiring a rethinking of organizational structures, cultural norms, and workflow integration (RQ3).
This study makes three key contributions to HCI and CSCW community:
  • First, we provide empirical evidence showing that GenAI adoption in newsrooms is currently individualized rather than collaborative. We highlight the gap between personal experimentation with GenAI usage and official organizational integration in collaborative work settings.
  • Second, we identify specific challenges that hinder GenAI integrations in newswork collaborations. We demonstrate how the absence of teamwork norms, coordination mechanisms, and shared practices impedes the emergence of accountable AI-augmented collaborative newsroom routines.
  • Third, we outline design and policy implications to inform the development of GenAI systems and organizational practices, fostering more effective integration of GenAI into collaborative work.
2. Related Work
We firstly review research on GenAI in the news industry (subsection 2.1). We then turn to broader studies of collaborations in the newsroom (subsection 2.2). Finally, we examine work on GenAI for teamwork and organizational collaboration (subsection 2.3).
2.1. GenAI in the News Industry
This subsection reviews prior work on GenAI in journalism, highlighting both its opportunities and risks (subsubsection 2.1.1) as well as the value tensions that potentially shape its cautious adoption in collaborative newsroom routines (subsubsection 2.1.2).
2.1.1. Opportunities and Concerns of GenAI in Journalism
Many news organizations now develop GenAI for drafting stories, summarizing reports, suggesting headlines, or even generating interview questions (Lewis et al., 2025; Ogola, 2024). Editing tasks are also supported through AI tools that propose alternative phrasings, check stylistic consistency, or flag potential bias (Noh et al., 2025; Wang et al., 2025). A 2023 global survey found that over 75% of news organizations across 46 countries were already using AI (JournalismAI, 2023). Some leading newsrooms such as China Media Group and Bloomberg are collaborating with AI technologists and experimenting with building tailored GenAI systems, seeking to customize GenAI for journalistic needs (Xiao et al., 2025b).
Despite these affordances, concerns persist. Chief among these are the risks of factual inaccuracy, the amplification of existing biases, and the homogenization of content, where widespread reliance on the same tools may diminish the distinctive voice and editorial identity of individual outlets (Siitonen et al., 2024; Lewis et al., 2025; Thomson et al., 2024; Matich et al., 2025). Journalists have also voiced existential anxieties about de-skilling and professional displacement, fearing that GenAI may automate precisely those creative and investigative tasks that define the integrity and social value of their work  (Woodruff et al., 2024; Siitonen et al., 2024). As Woodruff et al. argue  (Woodruff et al., 2024), these threats are particularly acute given the journalism industry’s already-precarious position, marked by economic instability, declining public trust, and shrinking editorial capacity.
HCI researchers have responded by exploring how GenAI tools can be aligned with newsroom practices (Aitamurto et al., 2019; Tseng et al., 2025; Xiao et al., 2025b; Nishal and Diakopoulos, 2025; Komatsu et al., 2020). For instance, Tseng et al.  (Tseng et al., 2025) proposed a journalist-centered design space for GenAI, arguing that systems must adapt to the editorial logics and workflows of the profession, rather than impose model-driven efficiencies. Hoque et al. (Hoque et al., 2024) designed a question-answering chatbot to assist journalists in presenting news content to readers in a balanced and objective manner.
2.1.2. Tensions Between GenAI and Newswork Routines
Several studies have highlighted how journalistic collective values such as independence, transparency, fairness, and credibility function not just as abstract ideals, but as design requirements for GenAI systems in newsrooms and sources of sociotechnical friction (Nishal and Diakopoulos, 2025; Møller, 2024; Komatsu et al., 2020; Xiao et al., 2025b). Nishal and Diakopoulos (Nishal and Diakopoulos, 2025) argue that journalistic values must be treated as central constraints in system development. Other scholars emphasizes that editorial judgment is a situated professional norm that must be maintained even as automation expands in newswork routines (Xiao et al., 2025b; Møller, 2024). Many newsrooms also formalized their AI policies (Hofeditz et al., 2025; Viner and Bateson, [n. d.]). Major organizations such as The Guardian have published explicit editorial guidelines on GenAI, focusing on the core values of the newswork when using GenAI (Viner and Bateson, [n. d.]).
These value tensions between GenAI and journalistic practice are not confined to top-down organizational directives. Recent studies show that journalists themselves are also experimenting with GenAI in informal, bottom-up ways (Wu, 2024). Wu (Wu, 2024) characterizes this as a “value-motivated use” framework, where journalists selectively adopt AI only insofar as it supports core editorial standards. In practice, individual journalists used GenAI for tasks such as designing interview questions, assisting with writing, clarifying terminology, summarizing key points, and checking facts. Yet these uses were marked by caution, as journalists had to constantly guard against potential errors, such as hallucinations, that could compromise their work (Wu, 2024). However, Wu (Wu, 2024) and other journalism studies do not examine how these individual practices relate to broader collaborative norms.
2.2. Collaborations in the Newsrooms
Newsrooms are fundamentally collaborative workplaces, where the production of news relies on distributed roles and interdependent tasks. Early ethnographic studies have emphasized that news is not authored by individual journalists alone but is the result of collective coordination (Tuchman, 1973; Cottle, 2007; Robinson, 2011). Collaboration structures how news stories are pitched, developed, verified, and ultimately published (Tuchman, 1973).
One core form of collaboration is the front-line journalist–editor relationship. Front-line journalists generate ideas, gather information, and draft stories, while editors provide oversight, frame narratives, and ensure adherence to professional values such as accuracy, fairness, and credibility (Frost, 2010; Kolodzy, 2006; Duffy, 2021). Duffy (Duffy, 2021) argued that much of journalism research tends to conflate the roles of front-line journalists and editors, overlooking the crucial distinctions between them. While front-line journalists are often foregrounded as the visible producers of news, editors play a central role as the gatekeepers of news values.
Another dimension of collaboration occurs among front-line journalists themselves. Carlson (Carlson, 2009) points out that front-line journalists, even when affiliated with different news organizations, often end up working in shared spaces such as press conferences, media briefings, or public events. They sometimes also directly coordinate with one another in joint reporting arrangements (Dailey et al., 2005; Carlson, 2009). The collaborations of front-line journalists create conditions where information is rapidly exchanged and professional norms are reinforced through peer interactions (Crouse, 2003).
Collaboration in newsrooms increasingly spans both internal professional boundaries and external organizational ones. Internally, front-line journalists and editors work closely with photojournalists, data journalists, designers, and technical staff to produce multimedia stories, navigating tensions between journalistic norms and technical or esthetic priorities (Xiao et al., 2025b; Wu et al., 2021; Pavlik, 2019). For instance, Xiao et al. (Xiao et al., 2025b) documented how in-house AI engineers and journalists collaborated to build GenAI tools: engineers providing technical infrastructure, and journalists contributing normative and contextual expertise. Externally, some today’s journalists also need to negotiate collaborations with technology companies that supply GenAI systems (Dodds et al., 2025; Xiao et al., 2025c; Simon, 2024); such partnerships then raise new questions about editorial independence (Simon, 2024).
2.3. GenAI for Teamwork and Organizational Collaboration
Before the rise of GenAI, a long line of research in HCI, CSCW and organizational studies examined how collaborative technologies, from email, shared document platforms, and groupware to chat tools, reshape workplace coordination and communication (Rama and Bishop, 2006; Rodden, 1991; Kong et al., 2024; Carasik and Grantham, 1988; Wang et al., 2022). These tools were primarily designed to improve information flow, enable distributed collaboration, and support decision-making across teams (Cherbonnier et al., 2025; Wang et al., 2022; Handel and Herbsleb, 2002). Early AI systems in organizations, such as expert systems or decision-support tools, were typically domain-specific and structured around narrow tasks (Billi and Labraña, 2025; El-Najdawi and Stylianou, 1993; Abdolmohammadi, 1987; Phillips-Wren, 2012). While useful, these systems rarely participated in open-ended dialogue or generative tasks. In contrast, GenAI introduces a new paradigm: tools that can autonomously generate, reframe, and negotiate meanings within collaborative work. This shift not only expands the scope of automation but also raises new questions about agency and the evolving role of human-AI collaboration (Leong et al., 2024; Shao et al., 2025; (2024), in partnership with Oracle).
According to Gartner’s 2024 survey (Gartner, 2024), GenAI has rapidly become the most widely adopted AI category in organizations. GenAI has opened up new opportunities for teamwork and organizational collaborations. At a basic level, GenAI is already widely used to support routine collaborative functions such as summarizing meeting notes, drafting shared documents, generating ideas, or providing quick answers and even feedback during group discussions. These functions help teams save time, reduce repetitive work, and establish a shared informational baseline (Palani and Ramos, 2024; Han et al., 2024; Takaffoli et al., 2024; Stray et al., 2025; Li et al., 2024a; Liang et al., 2024). More recently, however, research has pointed to emerging developments that extend GenAI’s role from a support tool to an active collaborator. For example, Leong (Leong et al., 2024) introduced the concept of Dittos, AI agents designed as personalized stand-ins that visually and aurally resemble humans and embody their knowledge. These clone agents were found to hold significant potential in team meetings, where they could evoke trust and a sense of presence while providing relevant information to support collective decision-making (Johnson et al., 2025).
Yet enthusiasm is tempered by significant barriers. Organizations struggle to demonstrate clear returns on AI, face shortages of skilled personnel, and encounter persistent data challenges such as access, privacy, and bias (Taherizadeh and Beaudry, 2023; Kleiman, 2025; Gartner, 2024). Organizational misalignment further complicates adoption: siloed initiatives and weak coordination hinder scaling, while knowledge management systems remain fragmented and difficult to build (Zhang et al., 2025b; Dodds et al., 2024). Concerns about trust, transparency, and legitimacy also create internal resistance among employees (Shao et al., 2025; Xiao et al., 2025b; Dodds et al., 2025).
3. Background: GenAI in the Chinese News Industry
In China, the news industry represents a distinctive setting due to its combination of rapid technological experimentation and strong regulation. The Chinese government has encouraged newsrooms to experiment with GenAI, including initiatives such as developing in-house GenAI systems through continuous recruitment of AI specialists and collaborations with technology companies and university labs (Xiao et al., 2025c; China Press, Publication, Radio, Film and Television Journal, 2024), reflecting a broader ambition to position China at the forefront of AI-driven media transformation. Meanwhile, Chinese newsrooms face significant accountability pressures: errors or misuses of GenAI can result not only in reputational damage among the public but also in institutional repercussions from government authorities (Xiao et al., 2025c). This provides an excellent case for examining how GenAI may potentially be integrated into collaborative newsroom routines.
The Chinese news industry has made notable progress in the design of the GenAI system for news work. [^1] Yet, it is worth noting that the development of these newsroom-specific GenAI systems has not been particularly transparent. For example, most Chinese newsrooms have not disclosed whether their “proprietary” models are entirely self-developed or fine-tuned from existing commercial or open-source foundations. All descriptions summarized here in the background section are directly from the state news reports about Chinese GenAI development in newsrooms, which limits independent verification of the underlying technical processes in this paper. At the organizational level, major media organizations have launched pilot projects that incorporate GenAI into routine news tasks. In July 2023, Xinhua News Agency (China Press, Publication, Radio, Film and Television Journal, 2024) released its proprietary large language model. The model was soon applied in coverage of the Chengdu Universiade, where Xinhua combined GenAI-based image and text generation to create and publish AIGC news videos (China Press, Publication, Radio, Film and Television Journal, 2024). Around the same time, China Media Group (CMG) introduced the “CMG Media GPT,” the first large-scale AI model in China dedicated specifically to audiovisual content production (China Press, Publication, Radio, Film and Television Journal, 2024; Lang, 2024). Also in February 2024, Jiefang Daily unveiled its own GenAI-powered application that could automatically reference the newspaper’s editorial database, and then produces articles in its distinctive newspaper style (China Press, Publication, Radio, Film and Television Journal, 2024). Even regional newspapers also accelerated their investment in domain-specific GenAI models. Zhejiang Radio and Television Group launched an AIGC-based platform to support professional news production and used it to publish news documentaries (China Press, Publication, Radio, Film and Television Journal, 2024). Jiangxi Daily has also developed a GenAI platform and could generate digital AI journalists for reporting (Feng, 2024).
These developments have drawn considerable attention in China’s media policy discourse. As noted by the China Press, Publication, Radio, Film and Television Journal (China Press, Publication, Radio, Film and Television Journal, 2024), which reflects the government’s stance and is directly overseen by the State Administration, such GenAI experiments highlight the need for journalists to develop new paradigms of human–AI collaboration in their daily newswork. Meanwhile, regulatory oversight in China has intensified. From January 2020, producing deceptive or misleading content with AI was made a criminal offense, aimed at deterring unethical applications of the technology (Xiao et al., 2025b). In 2023 the authorities released new rules governing the use of AI in news, which combined stricter accountability with incentives for news organizations to expand the use of GenAI in content production (of China, 2023). Starting from September 1, 2025, the Chinese government requires that all AI-generated content be clearly labeled. Any organization or individual found to maliciously delete, alter, forge, or conceal such labels will be considered in violation of the regulation and subject to penalties (CCTV News, 2025). It is worth noting that the development and application of GenAI in the news industry reflects broader state ambitions to position AI as a driver of technological leadership and cultural influence (Ding, 2018; Zeng, 2020).
4. Method
To investigate how GenAI is being integrated into collaborative newsroom routines, we conducted 27 in-depth interviews with senior newsroom managers, editors, and front-line journalists in China. All participants were affiliated with news organizations that had initiated early-stage projects involving the development or deployment of GenAI tools. Each journalist had direct experience with GenAI in their collaborative newswork. In the following, we outline our participant recruitment strategy, the design and conduct of the interviews, and our approach to data analysis.
4.1. Participants
We used purposive sampling (Firdaus et al., 2024) to recruit journalists from potential Chinese newsrooms actively engaged in early-stage GenAI experimentation or implementation. We ultimately interviewed 27 news practitioners (see Table 1). All participants were working in news organizations that had either (a) integrated general-purpose large language models into their workflows, (b) developed domain-specific GenAI models tailored for journalism (e.g., models fine-tuned on news language, ethics, and practices to support journalism broadly), or (c) implemented newsroom-specific GenAI tools designed for concrete editorial tasks such as article drafting, summarization, fact-checking, or text and video news production. A prerequisite for inclusion was direct experience with GenAI in daily newswork.
Table 1. Participant Demographics.
ID Gender Age Journalism Experience Position GenAI Experience
P1 Female 48 22 years Editor/Senior Manager 1, 2
P2 Female 42 19 years Editor/Senior Manager 1
P3 Male 45 23 years Editor/Senior Manager 1
P4 Female 38 13 years Editor 1
P5 Female 34 11 years Editor 1, 2
P6 Female 33 10 years Editor 2
P7 Male 37 10 years Editor 1, 2
P8 Male 35 12 years Editor 2
P9 Male 29 3 years Editor 2
P10 Male 28 3 years Front-line Journalist 2
P11 Female 26 2 years Front-line Journalist 2
P12 Female 24 2 years Front-line Journalist 2
P13 Male 32 8 years Editor/Front-line Journalist 2
P14 Female 30 6 years Editor/Front-line Journalist 2
P15 Female 28 3 years Front-line Journalist 1, 2
P16 Female 26 2 years Front-line Journalist 2
P17 Male 25 1 year Front-line Journalist 2
P18 Male 28 3 years Editor/Front-line Journalist 1, 2
P19 Male 25 1 year Front-line Journalist 2
P20 Male 25 1 year Front-line Journalist 2
P21 Male 25 1 year Front-line Journalist 2
P22 Male 25 2 years Front-line Journalist 2
P23 Male 26 1 year Front-line Journalist 2
P24 Female 26 2 years Front-line Journalist 2
P25 Female 26 1 year Front-line Journalist 2
P26 Female 25 1 year Front-line Journalist 2
P27 Male 25 1 year Front-line Journalist 2
Note: GenAI experience codes are defined as follows: 1 = Participation in organizational GenAI strategy or development planning; 2 = Use of GenAI tools in daily newswork routines.
To enrich our data with diverse perspectives and levels of organizational embedding, we sampled across different seniority levels, including editors, senior managers, and front-line journalists. Given our research goal of understanding how GenAI could be integrated into collaborative newsroom routines, we purposely recruited participants who had already engaged with GenAI in their professional roles. Our aim was to examine how journalists with firsthand experience actively negotiate GenAI’s role in teamwork and editorial coordination. While this strategy may bias the sample toward more AI-engaged participants, such a focus is appropriate for capturing the situated practices, tensions, and aspirations surrounding GenAI’s integration into collaborative newswork. Rather than studying those unfamiliar with GenAI and never thinking about future collaborative newswork around GenAI, we sought insight from those already reflecting on its collaborative implications.
17 participants were recruited through the authors’ personal networks, drawing on prior experience in journalism in China, while the remaining 10 were recruited through snowball sampling based on participants’ professional contacts. This approach allowed us to reach a relatively new and specialized population by leveraging trusted professional networks. These participants were not only early adopters of GenAI, but also embedded in organizational environments at the forefront of GenAI innovation. This recruitment strategy is especially appropriate for studying emerging practices that are not yet widely distributed or publicly documented, where recruitment through general channels would be difficult or ineffective (Dicce and Ewers, 2021).
Participants’ journalism experience ranged from one to over twenty years, and they represented a spectrum of positions including front-line journalists (n=16), editors (n=8), and senior managers in newsrooms (n=3). Some participants simultaneously held multiple roles within their organizations, as shown in Table 1. Front-line journalists offered grounded insights into how GenAI tools affect day-to-day reporting in collaborations. Mid-level editors, positioned between strategic directives and ground-level implementation, shared perspectives on coordinating GenAI integration into team workflows and balancing technological adoption with editorial standards. Senior managers contributed a broader organizational view, describing strategic rationales behind GenAI adoption, concerns about institutional credibility, and long-term visions for human–AI collaboration in journalism. Some editors also used GenAI to support their editorial tasks, and some front-line journalists occasionally participated in organizational discussions around GenAI strategy.
To protect participant confidentiality, all individuals were anonymized and assigned pseudonyms in the form of participant IDs ranging from P1 to P27. All participants provided informed consent prior to the interviews, and the study received ethics approval from our institutional review board.
4.2. Study Design
To explore how journalists experience, negotiate, and reflect on the integration of GenAI into their teamwork routines, we conducted 27 semi-structured interviews. Each interview lasted approximately 90–120 minutes and was conducted remotely. The study was approved by our university’s IRB, and all participants provided informed consent.
Each interview was structured into three phases. First, participants were asked to recall and describe concrete experiences with GenAI in their newsroom work, with particular attention to how these experiences unfolded in collaborative contexts. Participants reflected on how GenAI fit (or failed to fit) into broader editorial routines and teamwork. They were also asked to comment on their organization’s broader AI strategy, if any, and how early-stage GenAI practices were coordinated across the newsroom.
Second, the interview shifted to expectations. We asked participants to imagine what ideal GenAI integration would look like in collaborative newswork. Example questions included: “In your ideal scenario, how would GenAI assist not only you but also your team?” and “What would make a GenAI tool feel reliable in teamwork?”
Finally, participants were invited to reflect critically on the gap between their collaborative ideals and current realities. We asked them to identify what prevents GenAI from being integrated into team workflows, what forms of resistance or uncertainty arise when colleagues work with GenAI, and what organizational or technical changes might better support collective integration. Example questions included: “Why do you think GenAI has not yet become a regular part of your team’s workflow?” and “What organizational or technological changes would make GenAI more compatible with your newsroom’s collaborative practices?”
4.3. Data Analysis
We analyzed the data using reflexive thematic analysis (Braun and Clarke, 2019). This approach enabled us to identify recurring patterns across the dataset while remaining attentive to the contextual specificities of participants’ lived experiences. Consistent with many practices in qualitative HCI research (Chinh et al., 2019; Xiao et al., 2025b), we did not conduct statistical measures of inter-coder reliability. Our methodological stance is grounded in an interpretivist paradigm, which views knowledge as context-dependent and socially constructed. The goal was not to generate statistically generalizable results but to develop situated understandings that reflect participants’ perspectives (Morgan and Nica, 2020). All coding and theme development was conducted collaboratively by three researchers. Continuous discussions served to refine codes and themes, enrich interpretations, and ensure that our conclusions were jointly reasoned and methodologically sound. In the following section 5, section 6, and section 7, we outline the key themes that emerged from our analysis. These findings themes are also summarized in Figure 1 for reference.
Figure 1. Overview of the Findings Structure
5. Findings: Still Individual Use, but with Expectations for Integrating GenAI into Collaboration (RQ1)
Our findings reveal that although GenAI use is already widespread at the individual level, it is highly personalized, informal, and largely invisible to colleagues (subsection 5.1). Yet, these individualized practices have not yet coalesced into team-based routines or newsroom-wide standards for future collaborations. Even so, many journalists described actively working to ensure that their own GenAI use, and that of their teammates, remained aligned with professional values in collaborative work by individual efforts. Despite the lack of formal coordination, journalists nonetheless express a strong desire for GenAI to become a part of their collaborative routines (subsection 5.2).
5.1. Widespread but Individualized Use of GenAI
Across our interviews, participants expressed a wide spectrum of attitudes toward the use of GenAI in the newsroom, ranging from pessimistic to optimistic. Some viewed GenAI as a potential threat to journalistic rigor, fearing that it might “destroy the entire editorial process, make journalists lazier and less reflective, and ultimately lead them to rely on GenAI-generated text at the expense of their own writing skills” (P22). Others held a more hopeful outlook, believing that “GenAI can significantly improve journalistic productivity by saving time on tedious tasks and freeing up energy for meaningful investigative work” (P25).
Despite these differences, there was broad consensus on one point: the use of GenAI by individual journalists is already widespread and likely irreversible. “It’s nearly impossible to stop individual journalists from using GenAI,” P15 noted. “The best we can do is figure out how to manage it in the organizational level.” P15 further contextualized this trend by drawing an analogy to broader societal shifts. “It’s not just journalists, everyone is using some form of GenAI nowadays,” P15 explained. “Even eight-year-old child today is using GenAI. If school teachers cannot stop the child from using it, how could we possibly expect to stop a professional adult journalist?”
P2, a senior manager in her newsroom who was responsible for major organizational decisions and frequent communication with government authorities, cited a 2024 report by the research team from Xinhua News Agency (Xinhua Research Institute, 2024), the China’s most prominent state newspaper, stating that “very few newsrooms have expressed total rejection of GenAI in organizational level. Most are simply waiting for the right moment, for the GenAI technology to mature enough to be formally introduced into collaborative newswork.”
In P2’s view, whether GenAI should eventually be formally integrated into newsroom routines is hardly a matter for debate, since the “widespread individualized use of GenAI” by journalists has already made its presence in newswork effectively irreversible; the real question is how this integration ought to take place. She emphasized that the issue was not contingent on any individual journalist’s pessimism or outright rejection of GenAI, nor was it realistically possible for newsrooms either to prohibit individual journalists from using GenAI or to reject the technology entirely at the organizational level. After all, P2 believed, individual journalists could always access GenAI on their personal phones or computers without informing the organization. As long as a subset of journalists within the organization were already making use of GenAI in their everyday work, skepticism at either the personal or organizational level could not meaningfully halt its diffusion.
It is also worth noting that in our study, all direct interactions between journalists and GenAI took the form of conversational exchanges with chat-based GenAI systems such as Qwen or DeepSeek. Participants typically engaged GenAI by typing prompts in natural language, requesting draft texts, summaries, translations, or stylistic adjustments. In other words, GenAI was used as an interactive writing partner rather than through more technical means such as API calls, system integration, or backend customization. This use of GenAI illustrates its deep overlap with the core tasks of news production, integrating directly into the writing and editing, while also reinforcing its status as a largely private, individualized practice that can be easily concealed from colleagues and editors.
5.2. Motivations for GenAI Integrations into Collaborative Newsroom Routines
In our sampled newsrooms, maintaining the alignment between GenAI usage and journalistic expectations currently depends almost entirely on individual effort. In newsroom practice, front-line journalists typically draft stories, exchange leads with peers, and submit copy to editors, who then review, revise, and make publication decisions. Within this layered process, front-line journalists who experiment with GenAI must not only ensure the appropriateness of their own outputs but also anticipate how these outputs will intersect with editors’ judgments and colleagues’ contributions. Most participants described the extra effort this entailed: double-checking GenAI-generated text before passing it along, quietly monitoring whether teammates might have relied on GenAI, and compensating for possible errors to avoid undermining collective products. Editors, too, were not merely gatekeepers reacting to GenAI-produced content; some such as P5 and P7 also adopted GenAI tools to support their own editorial work, for example, using them to generate alternative headlines. These editorial uses similarly required individualized scrutiny to ensure alignment with house style and professional standards.
Yet such vigilance often left participants feeling isolated and burdened. Despite journalists’ increasing comfort using GenAI on their own terms, many participants expressed uncertainty about how their use intersected with that of their colleagues. Few had insight into how others were deploying GenAI in daily work, and most described their usage as “private” (e.g., P9, P13, P14) or “informal” (e.g., P5, P7, P16), a tool kept largely to themselves rather than openly discussed or coordinated within teams.
This dynamic not only reveals the professionalism and ethical awareness of individual journalists, but also underscores an urgent need for collective mechanisms that can support the responsible use of GenAI. Several participants, including P2, P4, P6, P8, P14, and P22, expressed deep concerns that the lack of formal integration of GenAI into newsroom workflows amplifies risks in collaborative work. Without shared guardrails or institutional guidance, they feared that GenAI could introduce errors into collective news products and causes team or even organizational level harms.
At the same time, participants saw promise in the idea that integrating GenAI into collaborative routines could help distribute responsibility more equitably across teams and reduce the cognitive and moral load currently borne by individual journalists. P12, for instance, explicitly questioned why no newsroom-wide policies or tools had been introduced, suggesting that management’s delay in institutionalizing GenAI might be a strategic way to offload risk onto individuals while avoiding organizational accountability.
Many participants (e.g., P3, P12 and P20) expressed a strong desire for more open conversations about the future role of GenAI in collaborative newswork. Rather than leaving the GenAI tool in a gray zone of informal, backstage experimentation for individuals, journalists hoped to see it open to scrutiny. They envisioned shared discussions about best collaborative practices, institutional expectations, and ethical tradeoffs as necessary foundations for building trust, enabling peer learning, and clarifying the legitimate place of GenAI in teamwork.
Notably, participants frequently contrasted GenAI with earlier personal productivity tools such as Microsoft Word or Excel. While these tools are similarly used by individual journalists, they were seen as passive instruments for documentation or formatting, with “no agency in shaping the content itself” (P7). GenAI, by contrast, was understood to be an active agent, generating, suggesting, and transforming journalistic text. Unlike word processors that could be mostly individualized, GenAI must be understood, coordinated, and made transparent within the newsroom collective.
In short, moving GenAI from the private domain into the realm of collective routines emerged as a widely shared aspiration across the newsroom contexts we studied.
6. Findings: Challenges in Integrating GenAI into Collaborative Newsroom Routines (RQ2)
In the section 5, we demonstrated that the current use of GenAI in newsrooms is highly individualized. Journalists engage with GenAI tools in personal, often private ways, with limited visibility into how their colleagues are using similar technologies. This fragmented usage stands in sharp contrast to the collaborative nature of journalistic work, which typically relies on shared routines, editorial review, and team-based production. In this section, we investigate why establishing collaborative GenAI routines has proven especially difficult in newsroom settings. It is important to clarify that our findings focus only on traditional newswork and do not extend to journalist-technical companies or journalists-AI engineers partnerships. This is because the technical teams or companies building GenAI tools do not participate in the day-to-day newswork.
These challenges in integrating GenAI into collaborative newsroom routines can be broadly grouped into two interrelated domains: structural and cultural. Structurally, many newsrooms lack formal coordination mechanisms for GenAI usage, making it difficult to establish shared visibility or accountability across team members (subsection 6.1). Culturally, journalists often regard their GenAI use as private, experimental, or even stigmatized, leading to reluctance in disclosing their practices or outputs to peers and supervisors (subsection 6.2).
6.1. Structural Challenges: Lack of Organizational Mechanisms for Integration
While individual journalists have actively experimented with GenAI tools, efforts to embed these practices into newsroom-level collaborative routines face critical structural barriers. These structural challenges stem from a lack of organizational incentives or mandates to coordinate GenAI usage (subsubsection 6.1.1), insufficient technical infrastructure to support collaborative workflows (subsubsection 6.1.2), and heightened public scrutiny that discourages formal integration without robust safeguards (subsubsection 6.1.3).
6.1.1. First, news organizations often lack structural incentives or mandates to coordinate GenAI usage in teamwork.
Although many newsroom leaders (e.g., P1, P2, P3) are aware that individual journalists within their organizations are already experimenting with GenAI tools, they rarely translate this awareness into concrete action at the team level. “We’ve done a lot to build internal GenAI systems, which is mainly to satisfy government expectations,” P13 explained, “but unfortunately, we don’t yet have a coherent collaboration guide about how to use GenAI for our journalists. No one really knows how it could enhance teamwork.” From a broader organizational perspective, P1, a senior manager, articulated this divide more explicitly. P1 emphasized that newsroom-level initiatives, often launched in response to national calls, tend to prioritize the development of specific GenAI systems, rather than paying close attention to how journalists incorporate GenAI into their daily practices.
What participants described as “GenAI systems” typically referred to newsroom-specific projects that fine-tuned existing large language models or built customized front-end interfaces to showcase AI-powered news functions. These prototypes, ranging from tools that automatically drafted short news items, to those generating video subtitles, summarizing interview transcripts, or retrieving archival content, functioned less as integrated aids to newsroom collaboration and more as deliverables that could be presented to authorities or internal management as evidence of technological advancement.
This gap between symbolic system-building and actual newsroom routines reflects a deeper uncertainty about how GenAI should be integrated into the labor structures of journalism. Several editors (n=7) voiced explicit concerns that formalizing GenAI use might unintentionally signal an endorsement of automation, thereby threatening long-standing norms around authorship, creativity, and editorial authority. P13 expressed this dilemma: “If we say it’s okay to use GenAI, would that encourage journalists to rely on it too much, instead of their own judgment? And how would our audience perceive this? Maybe they’d think, ‘Oh, this newspaper is just letting AI write the news now,’ or ‘Their journalism must be so poor that AI can step in.’ Some news audiences still carry a lot of negative impressions about GenAI.” This institutional hesitation creates a structural vacuum: GenAI becomes a tolerated but unregulated practice, operating in a gray zone between tacit acceptance and formal prohibition.
Moreover, even in organizations that recognize the potential value of collaborative GenAI use, many newsrooms lack the expertise or confidence to define what such collaboration should look like. As senior manager P3 put it bluntly: “We don’t even know how to try. What does GenAI collaboration actually mean?” P3 contrasted this ambiguity in collaboration with the current state of highly individualized GenAI use, where accountability remains clear. “Look, if everyone is using GenAI in a personal and unofficial way, and something goes wrong, we know who to hold responsible: it’s the individual journalist, right?” However, P3 raised concerns about what might happen if the newsroom were to formalize guidelines for how GenAI should be integrated into team workflows. “Now imagine we create an official framework for collaboration between human journalists and GenAI. Let’s say we even publish a clear policy about how GenAI shall be incorporated into teamwork. Then if a problem arises, who’s accountable? The journalist might say, ‘I was just following the policy.’ So where does that leave us?”
P3’s observation points to a critical tension: any move to institutionalize GenAI collaboration may inadvertently blur lines of responsibility, making it harder for editors or managers to attribute fault when mistakes occur. In contrast, the current hands-off approach, where GenAI use is quietly tolerated but not officially endorsed, preserves individual accountability and limits organizational liability. This ambiguity creates a paradoxical situation. On the one hand, organizations hesitate to create collaborative structures for fear of enabling blame deflection and institutional risk. On the other hand, without formal coordination mechanisms, GenAI use remains fragmented and inconsistent, inhibiting the development of scalable, transparent, and trustworthy practices. As a result, many newsrooms remain trapped in a structural limbo, aware of GenAI’s potential to reshape workflows, yet unwilling to define what responsible collaboration actually entails.
6.1.2. Second, the technical infrastructure of most Chinese newsrooms poses significant challenges to building collaborative GenAI routines.
Most GenAI tools were accessed individually via external platforms, such as Qwen Chat, Doubao, or DeepSeek. There are no built-in mechanisms in newsrooms to track when, where, or how GenAI has been used in the production of a specific story. Editors and collaborators often lack visibility into which portions of a draft were generated by AI, making it more difficult to assess the originality or editorial appropriateness of the content. As P2 explained, “People are mostly using GenAI on external platforms, which makes it hard for us to trace how they’re using it. We’re trying to build in-house GenAI systems with technical companies, but those are still in pilot stages. In most cases, journalists are still using external GenAI tools.”
In addition to traceability issues, GenAI tools lack shared interfaces that support collaborative review and editing. Unlike tools such as Google Docs, which facilitate real-time co-editing and inline commenting, most GenAI platforms are designed for single-user interaction. Journalists often copy and paste GenAI outputs into shared documents or use them as prompts for team discussions without revealing the prompts they used, the iterations they tried, or any potential misunderstandings by the model. These limitations are compounded by the fact that GenAI is fundamentally a black-box system. Editors, especially senior editors without a technical background, often lack the AI literacy needed to critically evaluate GenAI outputs. As P3 remarked, “On one hand, we don’t know how the journalist interacted with the GenAI. On the other hand, we don’t know how the GenAI came up with its response. It’s like two black boxes stacked on top of each other, which creates real challenges for editors.” This dual opacity of both human prompting and machine output undermines editorial confidence and creates a major barrier to integrating GenAI into collaborative newsroom routines.
The organizational structure of most newsroom tech teams further exacerbates these technical constraints. In our interviews, P1, P2, and P3 all acknowledged that their teams of developers were typically small, under-resourced, and primarily focused on building core GenAI functionalities, such as fine-tuning models or ensuring data compliance, rather than designing interfaces that support multi-user collaboration. As P1 explained, “Right now we’re still at the stage of putting food on the table, not worrying about how good the meal is. We need to get a basic GenAI platform running internally in the newsroom first to meet Chinese government’s expectations.” P1 further mentioned, “Only then can we think about building collaborative platforms. So for now, we let journalists explore freely and use whichever GenAI tools they prefer, instead of developing a system for coordinated use. We just don’t have the capacity.” This prioritization of backend capabilities over collaborative usability limits the newsroom’s ability to develop traceable and shared GenAI workflows.
6.1.3. Third, as nearly all editor participants in our study emphasized, reviewing GenAI-generated content and anticipating its potential risks has become increasingly difficult, especially as public awareness and sensitivity toward AI-generated news continues to grow.
This enhances the challenges for front-line journalist-editors collaborations around GenAI and the hesitations to promote an official guideline to integrate GenAI into newswork collaborations. Editors are no longer only concerned with factual accuracy or stylistic quality; they must also anticipate how GenAI usage will be perceived by increasingly critical audiences.
P14, for example, regularly monitors discussions on Chinese social media platforms such as Xiaohongshu and Weibo. She observed that readers are becoming more vigilant and suspicious about the presence of GenAI in journalistic content. As she described, “In the past, it was enough to ensure that the content met quality standards. Audiences wouldn’t notice or care whether AI was involved. But this year, especially in 2025, that has changed. You’ll find a lot of social media users actively trying to identify whether certain news articles were written by GenAI. And if they think the content was AI-generated, they feel disappointed. Sometimes these posts go viral, with thousands of likes, shares, and comments.”
This shift in public scrutiny introduces reputational risks for newsrooms and adds further pressure on editorial teams to detect GenAI usage, that are already structurally and technically difficult to manage, as discussed earlier. Without systematic mechanisms for tracing GenAI contributions or standard protocols for disclosure, editors face heightened audience expectations with limited organizational support.
As a result, newsroom leaders such as P1 and P2 are particularly cautious about openly integrating GenAI into editorial routines. Many fear that any premature or poorly safeguarded disclosure of GenAI use could trigger criticism, undermine credibility, and damage institutional trust. This apprehension leads organizations to delay or avoid formalizing GenAI practices, as external scrutiny amplifies internal hesitation and compounds the difficulty of embedding GenAI in collaborative workflows.
6.2. Cultural Challenges: Rethinking Journalists’ Identities in GenAI-Augmented Collaborative Work
Culturally, journalists often perceive the use of GenAI as private, experimental, or even stigmatized, leading to reluctance in disclosing their practices or outputs to journalist peers and editorial supervisors. This secrecy poses a major obstacle to integrating GenAI into collaborative newsroom workflows. Many journalists do not yet regard GenAI as a legitimate component of professional practice, but rather as a gray-area personal assistant. Because these practices are seen as outside the boundaries of “real journalism,” they are hidden from view, fragmenting knowledge and preventing the development of shared norms.
P24 explicitly described using GenAI as a form of professional shame, likening it to “cutting corners when no one is watching.” She explained that even if the final output is polished through human editing, the mere fact that the first draft was machine-generated made her feel like she had “cheated on an exam.” For her, GenAI was not a neutral tool, but a shortcut that undermined the pride she took in crafting language herself. “It’s like copying someone else’s homework and rewriting it to make it sound like your own,” she added. “Even if no one finds out, you still know what you did.”
Other metaphors further reinforced this sense of concealment. P27 referred to using GenAI as “ghostwriting.” She remarked, “Ghostwriting is not uncommon in journalism, especially in commercial collaborations. In many Chinese news reports involving corporate clients, the company drafts the content and the newspaper simply publishes it under a journalist’s byline. But it’s never openly discussed. It’s a hidden shame. If we don’t talk about ghostwriting, why would we publicly talk about GenAI?” In this framing, GenAI’s most powerful capability—text generation—is morally likened to unethical journalistic practices such as plagiarism or covert authorship and is not celebrated as innovation. These comparisons reflect not only individual discomfort, but also the normative environment in which journalists operate: one in which journalistic value is rooted in human creativity, intellectual labor, originality, and textual craftsmanship.
Many participants expressed concern that GenAI could devalue the profession itself, transforming journalism from a human enterprise into something replaceable by machines. As P24 self-mockingly noted, “We often talk about what professions AI might replace. News reports usually say it will replace low-creativity jobs like data entry. But if even high-literacy, high-creativity journalism can be assisted by GenAI, doesn’t that say something about how little society values our work?” T2 shared a widely discussed case in the Chinese news industry that illustrates this anxiety. T2 said a prominent Chinese newsroom recently sought to develop its own GenAI system for news writing and asked journalists to contribute training data by submitting ten of their most accomplished past articles. Yet, almost no journalist complied. As T2 explained, journalists feared that handing over their best work would not only strip them of control over the copyright of their writing but also allow the system to absorb their style and voice, erasing the individuality of their labor. For many, the idea that their copyrighted articles could be repurposed as raw material for a machine epitomized the risk of being replaced by the very system trained on their work. This moral framing of authorship, as something earned through struggle and talent, helps explain why GenAI remains a taboo topic, even within news organizations that encourage experimentation.
Fear of peer and supervisory judgment also played a critical role. In environments where editorial authority is tied to experience and skill, admitting to AI assistance might be interpreted as a lack of competence or diligence. Junior journalists especially felt that revealing GenAI usage might be seen as laziness or unprofessionalism. P27 reflected, “My editor is very old-school. If I told him I used AI to structure a piece or just organize interview notes, he’d think I wasn’t taking the job seriously.” This perception of stigma reinforces secrecy and prevents GenAI from being openly incorporated into collaborative work, even when it could enhance efficiency or creativity.
Additionally, the experimental nature of GenAI tools exacerbates their cultural marginalization. Many journalists are still figuring out how to use these tools effectively, and are therefore reluctant to share incomplete or uncertain practices. P25 expressed embarrassment about this phase of exploration: “I don’t know if other journalists are using GenAI. My computer skills are very low. I’m just playing around.” This hesitation further isolates GenAI use and prevents organizations from building collective intelligence around GenAI integration.
7. Findings: Opportunities Toward Future GenAI-Augmented Newswork Collaboration (RQ3)
Our participants further articulated concrete visions for how GenAI could be embedded into the future of collaborative newswork. Our findings show that journalists anticipate GenAI-augmented collaboration will require rethinking organizational priorities (subsection 7.1), fostering a culture of responsible use (subsection 7.2), and developing better infrastructures (subsection 7.3) and new routines (subsection 7.4) that embed GenAI into collaborative practice.
7.1. Rethinking Organizational Prioritization: Making Collaboration a Core Design Goal
Toward the end of many interviews, participants began to reflect critically on the organizational priorities that have shaped their newsroom’s GenAI projects. A recurring concern was the observation that internal AI teams and technical leadership tend to prioritize model development and task-specific performance over the collaborative workflows and user interfaces necessary for these tools to function meaningfully in daily editorial practice.
As P1 explained, “When resources are limited and especially when the government wants to see concrete GenAI system design outcomes, such as a functional prototype system that can be showcased in public or to the government as evidence of progress.” This mindset, while understandable in a results-driven institutional context like China’s, implicitly relegates collaboration to a downstream concern: something to be added later, once the core AI infrastructure is in place, rather than a foundational element of system design.
To address this imbalance, participants called for a cultural shift in GenAI development: collaboration must be treated as a first-order concern, not an afterthought. They outlined three specific strategies for achieving this transformation: First, involving editors and journalists early in the development process. Rather than viewing users as passive recipients of prebuilt systems, participants argued that journalists must be embedded in the development cycle from the beginning to ensure GenAI tools are grounded in collaborative realities. As P18 put it, “If no one is thinking about how these tools work in collaboration, how our colleagues or editors might perceive their use, it’s hard to trust that the tools are reliable in a newsroom collaboration context.”
Second, embed collaboration into the technical design of GenAI systems. Participants emphasized that GenAI platforms must support collaborative features, such as shared editing environments, real-time commenting, and visible histories of GenAI interactions. As P8 noted, “As an editor, I need to know how journalists interact with GenAI. If that visibility were built into the platform, it would help. But without it, we obviously can’t monitor every journalist’s chat history with GenAI, that would be outrageous and a violation of privacy.”
Third, expand GenAI evaluation metrics to include collaboration outcomes. Currently, internal GenAI pilot programs and evaluations often focus narrowly on performance: how well the system generates news text, images, or video. But participants argued that such metrics miss a critical dimension: whether GenAI improves or undermines collaborative newswork. As P12 suggested, “I’m not sure how to measure it, but maybe future evaluations should also include whether these tools actually enhance or hurt teamwork.”
These reflections point to a broader organizational insight for future GenAI design: its true value lies in how well it supports collaborative work in journalism: editorial judgment, team coordination, and collective accountability.
7.2. Fostering a Cultural Shift in Newsrooms: Embracing Responsible GenAI Use as Legitimate Newswork
Several participants called for a broader cultural transformation within their organizations: the normalization and legitimization of responsible GenAI use in routine journalistic practices. Although GenAI tools are already widely adopted by front-line journalists and editors individually, participants noted that their use remains largely hidden, ad hoc, or even stigmatized. Journalists worry that acknowledging reliance on GenAI might be interpreted as laziness, a lack of editorial rigor, or diminished professional credibility, especially in legacy newsrooms where traditional reporting values continue to dominate.
To address this hesitation, participants emphasized the importance of building a newsroom culture in which responsible GenAI use is not only accepted but also recognized as a legitimate form of editorial labor. P20 remarked, “We shouldn’t pretend everyone is doing everything by hand. That’s not honest, and it discourages people from trying. The reality is, many journalists are already using GenAI. We can’t pretend no one is.” P20 further explained that if GenAI use is becoming inevitable, then the next step must be to promote responsible use, not prohibit it, particularly given the practical impossibility of policing private GenAI use on personal devices.
To support this shift, participants urged newsroom leadership to provide clear internal messaging: the use of AI tools, when transparent, accountable, and thoughtfully integrated, should not be seen as undermining journalistic professionalism. In contrast, such use may reflect a strategic and informed approach to emerging technologies in pursuit of journalistic goals.
Participants also suggested that this cultural shift requires redefining what it means to be a responsible journalist in the GenAI era. Instead of viewing journalistic integrity as the absence of AI involvement, integrity should be evaluated based on the transparency and oversight exercised in using these tools. P3 emphasized, “Responsible GenAI use should be part of our professional standards, not something we hide from our peers or our audience.”
In participants’ vision, this transformation would also create space for open dialogue and peer learning around GenAI practices. “Journalists could more freely exchange prompt techniques, usage tips, and editorial guidelines” (P8), while “teams could collaboratively reflect on how GenAI use aligns with shared values” (P12).
Ultimately, legitimizing responsible GenAI use is a necessary step toward building a more adaptive, collaborative, and future-oriented newsroom culture.
7.3. Embedding GenAI into Newsroom Infrastructures for Future Collaboration
Participants emphasized the importance of integrating GenAI tools into newsroom technical infrastructures to support future collaborative work. Currently, most journalists rely on external interfaces, such as Qwen Chat, Doubao, or DeepSeek, to experiment with GenAI in their everyday newswork. This results in fragmented workflows and individualized practices that are invisible to other peers. Editors often cannot determine when GenAI was used, what prompts were given, or how much of the final draft was AI-generated, making it difficult to assess originality, editorial appropriateness, AI-like stylistic traces, or factual accuracy.
In contrast, according to our participants, embedding GenAI functionalities directly into newsrooms’ internal content management systems could potentially help transform GenAI from a siloed tool into an auditable and collectively accountable part of journalistic work. According to P3, if all interactions, such as prompts, generated drafts and post-editing revisions, were automatically captured and stored on a central platform of the newsroom, editors would be able to trace AI involvement, clarify accountability and make informed editorial decisions in the future. P3 put it: “We already have plans to build our own in-house GenAI system. What we need now is simply to integrate how people are already using GenAI into that system.” P1 agreed that, news organizations already investing in GenAI development can extend that investment to support collaborative visibility and editorial control. Such an integration also safeguards collaborative routines. It enables editorial teams to work from a common source of truth, reduces the burden on journalists to self-report AI use, and mitigates the reputational risks of AI-generated content slipping through unnoticed.
In this context, the integration of GenAI is also a way to maintain institutional legitimacy and prepare for potential government audits or public scrutiny regarding how AI is used in sensitive editorial decisions. In some state-run newsrooms, such as CCTV and CGTN, every piece of news must already pass through a centralized editorial platform that records each publication decision. As P1 explained, “We already have content management systems where every news article is tracked and every editorial decision is logged. So it makes sense to extend that to AI as well, especially in China, where we need to be prepared to justify our editorial decisions at any time, whether to the government or to the public.”
In P1’ imaginations, this kind of integration supports risk mitigation at scale. By embedding GenAI into existing editorial infrastructures, media organizations can respond more quickly to reputational, legal, or political concerns if AI-generated content is later questioned. It also creates a shared point of accountability for team members, strengthening collective responsibility and reducing the likelihood that GenAI use remains a hidden, individualized practice outside the editorial chain of command.
7.4. Creating New Work Routines around GenAI
Several participants proposed creating entirely new work routines that are uniquely enabled by GenAI, workflows that would be difficult, if not impossible, to realize under traditional human-only newswork. These routines do not simply replicate existing editorial practices with AI assistance; rather, they reflect a radical transformation in how content is produced, how collaboration unfolds, and how audiences engage with journalistic output in the age of GenAI.
Firstly, a common suggestion was the creation of dedicated AI news zones within newsrooms: editorial spaces deliberately separated from core news operations, designed to foster experimentation with GenAI tools. As P2 put it, “In these AI news zones, people know it’s okay to try something new. We’re not pretending it’s traditional reporting. It’s a different GenAI product.”
These AI zones were described as low-risk, high-flexibility environments where journalists teams could develop novel formats and experiment with creative applications of GenAI. Participants envisioned a wide range of possibilities that would be difficult to achieve through conventional human-only news production. For instance, some proposed GenAI-powered chatbots that “allow audiences to converse with digital journalist avatars to better understand complex news stories” (P22). Short-form videos and animations, scripted with GenAI and targeted at younger, mobile-first audiences, were also frequently mentioned, with P7 noting that “there’s already a lot on Douyin, and they’re popular”. Still others proposed AIGC documentaries composed from archival footage and GenAI-generated narration. P11 explained: “It’s technically feasible, and we have extensive archives, especially rich local footage that can be used to create documentaries about local cultures.”
Because these GenAI products are clearly distinct in form and tone from conventional journalism, our participants believe audiences tend to be more tolerant of stylistic experimentation and minor inaccuracies. Editors, in turn, feel freer to iterate on editorial norms. As P11 reflected, “When people see AI-generated content in these special AI news zones, they don’t judge it heavily by the standards of our front page. They expect it to be more innovative.”
Opening new routines for GenAI does not eliminate editorial control in our participants’ expectations. As P8 emphasized, even content produced in AI news zones must undergo human review, be clearly labeled as AI-generated, and be documented internally. However, the emergence of these formats offers a structured space to explore the creative frontier of GenAI without forcing immediate reform of core editorial systems. “In these zones,” P8 added, “journalists and editors can be more creative in how they collaborate with GenAI, without fear of professional judgment or reputational risk.”
Secondly, several participants also envisioned a new collaborative newswork form augmented by GenAI. P2 highlighted an experimental system called Fabarta Agents (E-commerce Research Center of China (2025), NetEase E-Commerce), currently being trialed at CCTV and several provincial news televisions as an internal newsroom application. A key feature of the GenAI system is its deliberate emphasis on peer learning and stylistic imitation. Journalists can recall the full corpus of a preferred journalist colleague’s past work in the system, using it not only as training material but also as real-time reference during writing. The agent can, for example, suggest opening paragraphs, headlines, or structural outlines based on how a trusted peer might have approached similar stories. As P2 explained, “You can ask the agent to write in the style of someone you respect in the newsroom. It’s like getting mentorship through the GenAI system.”
Importantly, Fabarta Agents was also designed to address copyright concerns that have accompanied other newsroom GenAI initiatives. According to P2, because the system operates as an internal application, journalists can upload their own past articles without fear that the content will be redistributed outside the organization or used for commercial purposes. In this way, copyright remains under newsroom control, and materials are treated as part of a closed professional resource, available only for internal reference or individual experimentation. P2 emphasized that this feature reassured many journalists who might otherwise resist contributing their work, as it positioned GenAI-assisted stylistic learning as a form of collegial exchange rather than a threat to ownership.
According to P2, Fabarta Agents opens up a new pathway for collegial influence and inspiration, where AI becomes a medium for professional emulation and reflection, not merely automation. As P2 summarized, “To me, this GenAI tool essentially creates a safe space for experimentation while also improving our collaborative culture. It shows how technology itself can reshape the very nature of collaboration.”
8. Discussion
Our findings suggest that while journalists widely adopt GenAI tools at the individual level, structural and cultural barriers prevent these practices from scaling into sustained team-based routines. At the same time, journalists envision productive futures for GenAI-augmented collaboration. These findings offer important design implications for HCI and CSCW researchers seeking to embed GenAI into collaborative work, particularly in high-stakes, institutionalized domains such as journalism.
This discussion section unpacks three interrelated dimensions. First, it calls for HCI and CSCW research to move beyond a predominant focus on individual interactions with GenAI and instead attend to its role in collaborative newsroom work (subsection 8.1). Second, it rethinks GenAI’s distinctive technical essence in contemporary collaboration, not as a peripheral support tool but as a system that directly intervenes in the epistemic and creative core of journalism, thereby reshaping how collaboration itself must be organized and legitimated (subsection 8.2). Third, it outlines concrete design and policy implications to support more transparent, accountable, and responsible forms of GenAI-augmented collaboration (subsection 8.3).
8.1. From Individual GenAI Use to Collaborative Work
While recent HCI and CSCW literature has explored the growing role of GenAI in creative and knowledge work, most of this research has focused on individual interactions with AI systems: how users prompt, evaluate, and iteratively shape GenAI outputs to support tasks such as writing, ideation, or coding (Palani and Ramos, 2024; Sun et al., 2024; Lee et al., 2025; Li et al., 2024b). In contrast, our study contributes to an emerging but underdeveloped area: how GenAI is (or is not) integrated into collaborative work practices, particularly in high-stakes, institutionally-bound domains such as journalism. It also raises forward-looking questions about how organizations can embed GenAI into future collaborative routines.
Our findings reveal that GenAI-augmented collaboration is not a natural extension of individual use, but a site of organizational tension, negotiation, and technical infrastructural fragility. Despite widespread individual experimentation with GenAI usage, newsrooms struggle to transform these fragmented practices into shared routines or team-level workflows. This aligns with long-standing HCI and CSCW insights that new technologies rarely slot in to existing work practices without reconfiguration of roles, norms, and communication structures (Møller, 2024; Holtgrewe, 2014; Blomberg and Karasti, 2013; Geiger et al., 2021; Hu et al., 2022). To some extent, adopting new technology also means adopting potential cultural or organizational transformations along with technology adoption. GenAI collaboration in our cases is a socio-technical challenge, encompassing both the design of collaborative interfaces and the need to ensure accountability, visibility, interpretability, and trust among newsroom actors.
In the technical aspect, our study suggests that the opacity of GenAI interactions, both in how prompts are crafted by journalists and in how outputs are generated by GenAI, renders shared understanding and joint decision-making difficult. This dual black-boxing of human prompting and machine generation creates asymmetries in editor–journalist collaboration, echoing prior HCI and CSCW concerns about collaboration breakdowns in common ground when collaborators cannot access or reconstruct one another’s interactions with complex systems (Carroll et al., 2006; Fischer, 2004; Ackerman et al., 2013). Editors, in our study, expressed a reluctance to incorporate GenAI into collaborative routines precisely because they lack technical mechanisms for auditability, revision tracking, or interpretive control.
Furthermore, from the cultural aspect, individual GenAI use tends to remain private, informal, and even stigmatized, which Goffman might call backstage behavior (Goffman, 1949). Goffman (Goffman, 1949) distinguishes between “frontstage” behavior, where individuals perform in alignment with institutional norms and expectations, and “backstage” behavior, which is concealed from public view and often deviates from formal roles. In our study, journalists’ use of GenAI largely occupies this backstage space: it is done alone, without disclosure, and often with ambivalence about its legitimacy. This concealment has important implications for how collaborative GenAI practices do or do not take shape. Because GenAI use is hidden, it resists normalization: it cannot be scrutinized, iterated upon, or collectively stabilized. As a result, organizations struggle to build legitimate, durable, and visible practices around AI collaboration.
While prior HCI studies have explored the technical design of AI tools to support individual workflows (Lu et al., 2025; Zhu et al., 2024), and CSCW scholars have examined the breakdown of collaboration in opaque systems (Easterbrook, 2012; Gui et al., 2018), our findings foreground a distinct dynamic: GenAI use becomes routinized individually without becoming institutionalized in collaborative work. This decoupling of individual technical use and limited social legitimacy, of what people do versus what they are willing to show, poses a unique barrier to scaling GenAI from individual to collaborative work.
While GenAI is increasingly adopted across real-world organizations, it has not yet emerged as an immediate or evident enabler of collaborative intelligence (Rajaram and Tinguely, 2024; Perifanou and Economides, 2025). This calls for further exploration into how GenAI can be more effectively integrated into organizational practices. In practice, moving from individual to collective GenAI use requires deep institutional scaffolding: shared norms about appropriate use, technical infrastructures that enable co-editing, and cultural shifts around disclosure and responsibility. We argue that GenAI does not “scale” to collaboration by default: it must be domesticated into organizational life through deliberate design and negotiation.
8.2. Rethinking GenAI as a Collaborative Technology in Future Newsroom
Unlike earlier collaborative technologies that operated behind the scenes, GenAI directly intervenes in the epistemic and creative core of journalism. This shift forces a renegotiation of routines, roles, and professional legitimacy. Prior HCI, CSCW and organizational research has shown how emerging technologies, such as groupware, workflow systems, and enterprise platforms, reshape collaboration by redistributing tasks, restructuring coordination, and altering authority (Rama and Bishop, 2006; Poltrock and Grudin, 1998; Grudin, 1999). These previous generations of tools, however, typically operated in the background: they supported communication, scheduling, or documentation, but rarely intervened in the substantive outputs of professional work (Rama and Bishop, 2006; Knearem et al., 2023; Xiao et al., 2025a). Consequently, their adoption was often gradual and partial, limited to situations where technology could be integrated without disrupting established routines of teamwork and accountability.
GenAI differs in a fundamental way. Rather than remaining in the background, it intervenes directly and radically in the epistemic and creative core of journalism by drafting headlines, summarizing interviews, or generating full article drafts, as we mentioned in subsection 6.2 and subsection 7.2. This direct involvement not only reshapes how work is done but also blurs professional boundaries of authorship, editorial authority, and responsibility. In contrast to earlier technologies that modified supporting processes, GenAI compels today’s journalists and editors to renegotiate what counts as legitimate professional newswork and who is accountable for it.
We argue that GenAI’s intervention in core editorial tasks fundamentally alters how collaboration unfolds in the newsroom. We found that journalists rarely disclosed their practices to colleagues or editors. This secrecy was driven not only by coordination difficulties, such as tracing AI contributions, but also by deeper anxieties. Many feared that reliance on GenAI would erode their professional skills, increase the risk of replacement, or compromise copyright and recognition of their work, echoing earlier HCI research on copyright concerns (Suresh et al., 2024) and secret GenAI use (Zhang et al., 2025a).
These findings highlight that, for GenAI to support rather than disrupt newsroom collaboration, it must move beyond hidden, individualized use, and become embedded in shared routines. This direct intervention from GenAI in the substance of journalistic work transforms collaboration at a fundamental level: questions of authorship, authority, and accountability are no longer peripheral adoption issues but central to everyday teamwork in the age of GenAI. Secretive reliance on GenAI, as discussed in subsection 5.2 and subsection 6.2, can by contrast actively undermine collaboration, eroding trust among colleagues and obscuring accountability for editorial decisions. These practices often stem from journalists’ desire to preserve their professional identity and safeguard the distinct value of human judgment, even as GenAI encroaches on the very tasks that define their role. Yet such consequences underscore the need to fundamentally reconsider GenAI as a technology that disrupts the epistemic and creative core of journalism, and to reimagine how collaborative practices can adapt in response.
We thus urge journalism scholars and practitioners to move beyond treating GenAI as a matter of private experimentation. Instead, attention should turn to how newsroom cultures can evolve to support more transparent, coordinated, and principled uses of GenAI that reinforce journalism’s collective standards of fairness, independence, and accuracy. Future HCI and CSCW research should investigate how GenAI’s integration into collaborative routines reshapes professional identity, and how, in this context, teams and organizations can renegotiate the boundaries of human judgment and GenAI contribution in sustaining the professional identity. More broadly, scholars must also consider how a technology as disruptive as GenAI can be embedded into future collaboration in ways that channel its interventions into collective, accountable, and sustainable practices rather than isolated or hidden uses.
8.3. Design and Policy Implication
We call for bringing GenAI use into newsrooms out of the shadows and embedding it into collaborative, transparent, and accountable practices that align with both everyday workflows and institutional values.
8.3.1. Designing for Collaborative GenAI Integration
First, design for social traceability to counter the “double black box.” In our study, GenAI use was often invisible on two levels: editors and colleagues rarely knew when or how journalists were using GenAI, and journalists themselves often lacked clarity on how the model produced its outputs. This double opacity not only undermines accountability but also weakens trust in collaborative processes. Design should therefore make both human and AI contributions more visible across the editorial workflow. For instance, interfaces could visually differentiate between human-generated and AI-generated text, display prompt histories, and show revision paths.
Second, design for shared editorial work flow. Current GenAI tools primarily operate in a single-user paradigm, where one journalist privately interacts with the system. Yet our participants increasingly envisioned GenAI as a shared infrastructure woven into team-based routines. To realize this vision, design should enable collaborative use: shared GenAI workspaces, prompt libraries editable by multiple users, and co-edit outputs that record collective contributions.
Third, turning institutional norms into interface-level affordance. Journalists sought tools that would help them uphold values of transparency, attribution, and editorial integrity. Systems can support this by prompting disclosure of GenAI involvement, suggesting attribution formats consistent with newsroom policies, or flagging content that might raise ethical concerns. Integrating these practices into the interface itself would allow journalists to enact shared values in everyday work, helping normalize GenAI as a legitimate, collectively governed part of newsroom routines.
8.3.2. Toward Collaborative and Accountable GenAI Policies
First, to meaningfully support GenAI-augmented collaboration, newsrooms must treat collaboration not as a downstream usability concern, but as a strategic design priority. Institutional policies should incentivize the early involvement of front-line journalists in AI development processes and require that GenAI evaluation metrics extend beyond technical benchmarks like accuracy. Instead, assessments should include collaborative usability, editorial interpretability, and shared accountability across team members.
Second, newsrooms must establish clearer norms and incentives to institutionalize the responsible use of GenAI. Our findings indicate that informal, often stigmatized use of GenAI persists in part because of the absence of formal guidance and supportive peer practices. Policies should not treat GenAI use as a compliance problem, but as a cultural opportunity to scaffold shared learning and responsible exploration. Practices such as shared prompt libraries, peer-learning forums, and transparent audit trails can foster a culture of open dialogue, where GenAI use is not hidden but collaboratively refined and improved.
Third, for GenAI to be governable and sustainable in the long term, it must be integrated into the newsroom’s editorial routines. We emphasize that GenAI adoption and integration must occur at the infrastructural level. Policies should require traceable logging of GenAI involvement in content production, guidelines for data retention, and clear delineation of editorial responsibility when AI-generated content is used. These mechanisms not only support internal accountability and collaboration, but also prepare news organizations to meet external demands, such as regulatory compliance, reputational risk management, and public scrutiny over AI usage in high-stakes reporting environments.
9. Limitations and Future Work
This study provides early insight into the emerging role of GenAI in newswork collaboration, but several limitations should be noted. First, our sample is limited to Chinese newsrooms, which operate within a distinct media system shaped by strong institutional norms and political constraints. While many of the structural and cultural dynamics we observed may resonate globally, future research should examine how GenAI adoption varies across different media ecosystems, including commercial, public service, and nonprofit news organizations in democratic contexts. Examining related sectors such as education, law, or healthcare could also offer valuable contrasts in how GenAI is adopted into collaborative practices in high-stakes domains.
Second, our recruitment strategy and study design may have favored journalists who used GenAI directly and were relatively open or positive toward GenAI. Because many participants were recruited through personal networks and snowball sampling, those who were skeptical, disengaged, or more deeply concerned about GenAI’s risks may be underrepresented. This limitation is compounded by the reliance on self-reported accounts, which can introduce biases in how participants frame their own practices and attitudes. Future work should seek to incorporate more diverse perspectives, including journalists who resist or avoid GenAI, and combine self-reports with direct observation or usage data to mitigate these biases.
Finally, our research focused primarily on newsroom staff, particularly editors and journalists. However, collaborative GenAI use may also involve non-editorial actors such as designers, developers, and product managers. Future work should adopt a more holistic organizational lens to trace how various teams negotiate GenAI integration.
10. Conclusion
Through 27 interviews with senior editors, newsroom managers and front-line journalists in China, this study examined how journalists navigate the integration of GenAI into collaborative newswork. We identified a clear gap between individual experimentation with GenAI and the structural and cultural mechanisms needed to support sustained, legitimate, and team-based use. While journalists expressed curiosity and even enthusiasm about GenAI’s potential, their current practices remain fragmented, private, and stigmatized. Our paper contributes to HCI and CSCW by foregrounding the sociotechnical barriers that prevent GenAI from becoming a collaborative infrastructure and pointing out future opportunities to integrate GenAI into collaborative work.